OpenStudy (anonymous):

How to rewrite the this as Ax+By+C=0 f(x)=5/2(2x+4)

OpenStudy (texaschic101):

f(x) is the same as y so we have : y = 5/2(2x + 4) -- distribute through the parenthesis y = 5x + 10 -- subtract 5x from both sides -5x + y = 10 -- multiply by -1 5x - y = -10 -- add 10 to both sides 5x - y + 10 = 0

OpenStudy (anonymous):

can you give the exact procedure i want to know how did u get that answer thnks

OpenStudy (anonymous):

there's a lot of it that i didnt understand

OpenStudy (anonymous):

your answer is not clear to me sir

OpenStudy (texaschic101):

f(x) represents y so your equation now is : y = 5/2(2x + 4) we now want to get rid of the parenthesis by multiplying the number outside of the parenthesis by everything in the parenthesis. So, basically, we are doing this... y = 5/2 * 2x + 5/2 * 4 y = 10/2x + 20/2 -- reduce y = 5x + 10 now we want this in Ax + By = C so first we can subtract 5x from both sides -5x + y = 5x - 5x + 10 -- combine like terms -5x + y = 10 now we can subtract 10 from both sides -5x + y - 10 = 10 - 10 (combine like terms) -5x + y - 10 = 0 we will now multiply the entire equation by -1, because we do not want x to be negative -5x + y - 10 = 0 ---->(-1) 5x - y + 10 = 0 <=== standard form does that help any ?

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

i get it now thnx

OpenStudy (texaschic101):

sure thing :)